# Random Utility with Aggregation

### v2

@@ 3.1 Vertex representation
In this section, we characterize admissible consumer behaviors under RU-rationality and show that many such behaviors are not ARU-rational. No assumption on the domain is required. However, to simplify exposition, we impose that every observable choice set contains the outside option $a_{0}$ (i.e., $a_{0}\in D$ for all $D\in\mathscr{D}$), consistent with the common view that a status quo alternative is always available. (Even without this assumption, the main message of the section remains unchanged; see Remark 3.11 for details.)
We first introduce a notation: for each $\succ\in\mathscr{L}(\mathscr{A})$ and $\mathscr{E}\subseteq\mathscr{D}$, define a function denoted by $c^{\succ}_{\mathscr{E}}$ over $\mathscr{D}$ as follows: for all $D\in\mathscr{D}$,
@@
$\rho^{\succ}_{\mathscr{E}}$ is RU-rational for all $\succ\in\mathscr{L}(\mathscr{A})$ and $\mathscr{E}\subseteq\mathscr{D}$. [^11] See Section A.2 in the appendix for the proof.
@@
Consider the outside option setup (i.e., $\mathscr{A}_{A}=\{a_{0}\}$). A stochastic choice function $\rho$ is RU-rational if and only if
@@
We refer to this set as the RU polytope. Moreover, each $\rho^{\succ}_{\mathscr{E}}$ is a vertex of the RU polytope.
Mathematically, the theorem provides a vertex characterization of RU-rational stochastic choice functions, whereas Theorem 3.1 offers a hyperplane characterization. The vertex characterization shows that RU-rationality permits a type of behavior excluded under ARU-rationality: the agent may sometimes choose the outside option (the default option) $a_{0}$, possibly due to menu effects, while in other cases adhering to a rational ranking.
To explicitly compare the RU polytope and the set of stochastic choice fictions that are ARU-rational, we first define the latter set as follows denoted by $\mathcal{P}_{\mathscr{A}}$:
@@
$\mathcal{P}_{\mathscr{A}}=\left\{\rho:\rho(A,a)=\sum_{\succ\in\mathscr{L}(\mathscr{A})}\mu_{{\scriptscriptstyle\mathscr{A}}}(\succ)\rho^{\succ}_{\mathscr{D}}(A,a)\quad\text{for all }A\in\mathscr{D},\ a\in A,\ \text{with }\mu_{{\scriptscriptstyle\mathscr{A}}}\in\Delta(\mathcal{L}(\mathscr{A}))\right\},$
where $\Delta(\mathcal{L}(\mathscr{A}))$ is the set of probability distributions over $\mathcal{L}(\mathscr{A})$ and $\rho^{\succ}_{\mathscr{D}}$ is defined by (6).
@@
The set $\mathcal{P}_{\mathscr{A}}$ is a polytope with vertices $\rho^{\succ}_{\mathscr{D}}$, or
$\mathcal{P}_{\mathscr{A}}=\text{co.}\{\rho^{\succ}_{\mathscr{D}}|\succ\in\mathscr{L}(\mathscr{A})\}.$
We refer to this set as the ARU polytope.
@@
Theorem 3.6 and Corollary 3.9 together show that the ARU polytope is strictly smaller than the RU polytope: it is a subpolytope whose vertices form only a small subset of the much larger collection of RU vertices. Specifically, the vertices of the RU polytope are given by $\rho^{\succ}_{\mathscr{E}}$ for all $\succ\in\mathscr{L}(\mathscr{A})$ and $\mathscr{E}\subseteq\mathscr{D}$, while the vertices of the ARU polytope are given by $\rho^{\succ}_{\mathscr{D}}$ for all $\succ\in\mathscr{L}(\mathscr{A})$. Thus, ARU vertices vary only with the ranking $\succ$, whereas RU vertices vary with both $\succ$ and the choice of $\mathscr{E}$. See Figure 4 in Section 5.3 for an illustration. The remark quantifies this gap:
@@
The number of vertices (i.e., the number of $\rho^{\succ}_{\mathscr{E}}$) of the RU polytope is double-exponentially larger than the number of vertices of the ARU polytope (i.e., the number of $\rho^{\succ}_{2^{\mathscr{A}}}$), provided that $\mathscr{D}$ is sufficiently rich. For example, if $\mathscr{D}$ contains all sets that include $a_{0}$, then the ratio of the number of vertices in the RU polytope to the number in the ARU polytope is at least
$\frac{2^{(2^{n}-{n\choose 2}-1)}}{n+1},$
where $|\mathscr{A}_{N}|=n$. Since the dominant term in the numerator is double exponential, the ratio grows extremely rapidly. In fact, when $n\geq 6$, the ratio already exceeds $2^{2^{\,n-1}}$.
Theorem 3.6 and Remark 3.8 and 3.10 together show that RU-rationality imposes far weaker restrictions than ARU-rationality, which is consistent with the main message of Theorem 3.1.
@@
As mentioned, the main result of this section (Theorem 3.6) does not require the domain assumption that $a_{0}\in D$ for all $D\in\mathscr{D}$. The theorem continues to hold with a slight modification of the definition of $c^{\succ}_{\mathscr{E}}$. The adjustment is needed because a set $D$ may fail to contain $a_{0}$ even when $D\notin\mathscr{E}$. The generalized definition is:
3.2 Multiple aggregated categories
In this section, we consider the case in which there exist multiple aggregated categories as well non-aggregated categories. The main results from the single aggregated category case continue to hold in essentially the same form, as we show below:
@@ 3.2 Multiple aggregated categories
A stochastic choice function $\rho$ is RU-rational if and only if $\rho$ satisfies the following two conditions:
@@
    (Limited Monotonicity): if $b\in D\subseteq\mathscr{A}_{N}$ and $E\subseteq\mathscr{A}_{A}$ then $\rho(D,b)\geq\rho(D\cup E,b)$,
@@
The only substantive difference from Theorem 3.1 is that the condition (i) becomes stronger. In the earlier theorem, the condition (i) required monotonicity with respect to the addition of a single aggregated category—namely, the outside option $a_{0}$. In the current theorem, where multiple aggregated categories are present, the condition (i) instead requires monotonicity with respect to the addition of arbitrary subsets $E$ of the aggregated categories. This is a natural extension of Theorem 3.1: Theorem 3.1 corresponds to the case in which $\mathscr{A}_{A}$ is singleton $\{a_{0}\}$. Therefore, all of the implications of Theorem  3.1, as explained in the previous section hold with Theorem 3.12.
The proof is constructive and more involved as we need to define a composition distribution, or the distribution function defined on the product space $\prod_{a\in\mathscr{A}_{A}}(2^{X(a)}\setminus\{\emptyset\})$, not just on $2^{X(a_{0})}\setminus\{\emptyset\}$. We provide the formal proof is in the appendix.
Remark 3.13.
In line with the single aggregated category case, the only domain assumption required for Theorem 3.12 to hold is that if $D\cup E\in\mathscr{D}$ for $D\subseteq\mathscr{A}_{N}$ and $E\subseteq\mathscr{A}_{A}$, then $D\in\mathscr{D}$. This assumption is only needed in order to define Limited Monotonicity and plays no conceptual role beyond that.
@@ 4 Conditions Restoring ARUM
This section investigates the conditions under which RU-rationality implies ARU-rationality. These results are especially relevant given the widespread use of ARUMs in empirical IO and the common assumption that the data-generating process satisfies RU-rationality. They provide concrete guidance on how to define aggregated categories in a way that justifies the use of ARUMs.
In particular, we provide sufficient restrictions on the underlying preference distribution $\mu_{{\scriptscriptstyle\mathscr{X}}}$ and the composition distribution $\lambda$ that ensure RU-rationality and ARU-rationality coincide on the observed choice domain $\mathscr{A}$. The two conditions we introduce are independent: one applies to $\mu_{{\scriptscriptstyle\mathscr{X}}}$ (Section 4.1) and the other to $\lambda$ (Section 4.2).
@@ 4.1 Condition on $\mu_{{\scriptscriptstyle\mathscr{X}}}$: non-overlappingness
In this section we provide a restriction on the distribution of preferences that recovers the implications of ARUM. The restriction says for each aggregated category $a$, the underlying alternatives in $X(a)$ are similar in terms of preference rankings.
@@
Given an aggregation function $X$, a distribution $\mu_{{\scriptscriptstyle\mathscr{X}}}\in\Delta(\mathscr{L}(\mathscr{X}))$ is non-overlapping if for any $\succ\in supp(\mu_{{\scriptscriptstyle\mathscr{X}}})$, $a\in\mathscr{A}_{A}$, and $y,z\in X(a)$,
@@
The definition says that any preference in the support of $\mu_{{\scriptscriptstyle\mathscr{X}}}$ has the property that each $X(a)$ does not overlap with other aggregated categories in terms of preference ranking. [^13] It is possible to weaken the definition as follows but we keep this condition for simplicity: A distribution $\mu_{{\scriptscriptstyle\mathscr{X}}}\in\Delta(\mathscr{L}(\mathscr{X}))$ is almost non-overlapping if for any $\succ\in supp(\mu_{{\scriptscriptstyle\mathscr{X}}})$, there exists $b,c\in\mathscr{A}$ such that (i) $y\succ x$ for all $x\in X(b)\cup X(c)$ and $y\in\mathscr{X}\setminus(X(b)\cup X(c))$ and (ii) for all $a\in\mathscr{A}_{A}\setminus\{b,c\}$, and $y,z\in X(a)$, $y\succ x\succ z\implies x\in X(a)$.
@@
Proposition 4.2 shows that, without any assumptions on $\lambda$, the use of ARUM is valid when the alternatives grouped within each aggregate occupy similar positions in agents’ preference rankings. Consequently, when constructing an aggregated data set, alternatives with closely related characteristics should not be assigned to different aggregates. A similar observation has been made empirically by Stafford (2018) using a dataset of lobster fishing. In particular, she demonstrated that classifying crab-fishing sites as part of the outside option leads to biased coefficient estimates. Because lobster and crab fishing are close substitutes, an excellent lobster site is preferred to a poor crab site and vice-versa, their rankings overlap, thereby violating the non-overlapping condition specified in Proposition 4.2.

### v3

@@ 3.1 Vertex representation
In this section, we provide a vertex characterization of admissible behavior under RU-rationality and show that many such behaviors are not ARU-rational. This analysis does not require any assumption on the choice domain $\mathscr{D}$. To streamline the exposition, however, we impose that every observable menu contains the outside option $a_{0}$ (i.e., $a_{0}\in D$ for all $D\in\mathscr{D}$), consistent with the common view that a default (status quo) alternative is always available. [^12] Even without this assumption, the main message of this section is unchanged; see Remark D.6 in the online appendix for details.
Recall that an ARU-rational stochastic choice function can be written as a convex combination of deterministic choice functions induced by linear orders over aggregates (see Remark 3.5 below). This means that ARU-admissible deterministic behaviors are precisely the rational choice functions generated by linear orders on the set of aggregates.
To characterize admissible consumer behaviors under RU-rationality, we first introduce the following notation: for each $\succ\in\mathscr{L}(\mathscr{A})$ and $\mathscr{E}\subseteq\mathscr{D}$, define a function denoted by $c^{\succ}_{\mathscr{E}}$ over $\mathscr{D}$ as follows: for all $D\in\mathscr{D}$,
@@
$\rho^{\succ}_{\mathscr{E}}$ is RU-rational for all $\succ\in\mathscr{L}(\mathscr{A})$ and $\mathscr{E}\subseteq\mathscr{D}$. [^13] See Section B.1 in the online appendix for the proof.
@@
Assume Richness of $\mathscr{X}$. Consider the outside option setup (i.e., $\mathscr{A}_{N}=\{a_{0}\}$). A stochastic choice function $\rho$ is RU-rational if and only if
@@
The set is called the RU polytope. Moreover, each $\rho^{\succ}_{\mathscr{E}}$ is a vertex of the RU polytope.
The proof of the theorem is in Section A.2 of the appendix. Mathematically, the theorem provides a vertex characterization of RU-rational stochastic choice functions, whereas Theorem 3.1 offers a hyperplane characterization. [^14] Partial-RU rationality can be characterized by linear inequalities–the non-negativity of BM polynomials when $\mathscr{D}=2^{\mathscr{A}}\setminus\{\emptyset\}$. The vertex characterization shows that RU-rationality permits a type of behavior excluded under ARU-rationality: the agent may sometimes choose the outside option (the default option) $a_{0}$, possibly due to menu effects, while in other cases adhering to a rational ranking.
To explicitly compare the RU polytope and the set of stochastic choice functions that are ARU-rational, we first define the latter set as follows denoted by $\mathcal{P}_{\mathscr{A}}$:
@@
$\mathcal{P}_{\mathscr{A}}=\{\rho:\rho=\sum_{\succ\in\mathscr{L}(\mathscr{A})}\mu_{{\scriptscriptstyle\mathscr{A}}}(\succ)\rho^{\succ}_{\mathscr{D}}\text{ for some }\mu_{{\scriptscriptstyle\mathscr{A}}}\in\Delta(\mathcal{L}(\mathscr{A}))\}$, where $\Delta(\mathcal{L}(\mathscr{A}))$ is the set of probability distributions over $\mathcal{L}(\mathscr{A})$ and $\rho^{\succ}_{\mathscr{D}}$ is defined by (5).
@@
The set $\mathcal{P}_{\mathscr{A}}$ is a polytope with vertices $\rho^{\succ}_{\mathscr{D}}$, or $\mathcal{P}_{\mathscr{A}}=\text{co.}\{\rho^{\succ}_{\mathscr{D}}|\succ\in\mathscr{L}(\mathscr{A})\}$. We refer to this set as the ARU polytope.
@@
Theorem 3.2 and Corollary 3.1 together show that the ARU polytope is strictly smaller than the RU polytope: it is a subpolytope whose vertices form only a small subset of the much larger collection of RU vertices. Specifically, the vertices of the RU polytope are given by $\rho^{\succ}_{\mathscr{E}}$ for all $\succ\in\mathscr{L}(\mathscr{A})$ and $\mathscr{E}\subseteq\mathscr{D}$, while the vertices of the ARU polytope are given by $\rho^{\succ}_{\mathscr{D}}$ for all $\succ\in\mathscr{L}(\mathscr{A})$ (i.e., $\mathscr{E}=\mathscr{D}$). Thus, ARU vertices vary only with the ranking $\succ$, whereas RU vertices vary with both $\succ$ and the choice of $\mathscr{E}$. See Figure 1 in the next section for an illustration. The next remark quantifies this gap:
@@
The number of vertices (i.e., the number of $\rho^{\succ}_{\mathscr{E}}$) of the RU polytope is double-exponentially larger than the number of vertices of the ARU polytope (i.e., the number of $\rho^{\succ}_{2^{\mathscr{A}}}$), provided that $\mathscr{D}$ is sufficiently rich. For example, if $\mathscr{D}$ contains all sets that include $a_{0}$, then the ratio of the number of vertices in the RU polytope to the number in the ARU polytope is at least $2^{(2^{n}-{n\choose 2}-1)}/(n+1)$, where $|\mathscr{A}_{A}|=n$. Since the dominant term in the numerator is double exponential, the ratio grows extremely rapidly. In fact, when $n\geq 6$, the ratio already exceeds $2^{2^{\,n-1}}$.
Theorem 3.2 and Remark 3.5 and 3.6 together show that RU-rationality imposes far weaker restrictions than ARU-rationality, which is consistent with the main message of Theorem 3.1.
3.2 Exogenously given aggregation correspondences
In the previous sections, we impose no restrictions on the aggregation correspondence $X(\cdot)$, reflecting the practical difficulty of enumerating all underlying alternatives represented by a non-atomic aggregate $a$. The framework, however, extends directly to the case in which the aggregation correspondence is fixed exogenously. This is relevant when the analyst knows the possible underling alternatives that each aggregates may represent but does not observe which of them are available to consumers in each market. For example, in a study of demand for specific automobile models, the outside option—not purchasing a car—may encompass known transportation alternatives such as public transit, cycling, and ride-sharing. The analyst can enumerate these alternatives, yet their availability varies across markets: some cities lack a subway system, others have no bike-sharing program.
For simplicity, we focus on the case in which there is only one non-atomic aggregate (i.e., $\mathscr{A}_{N}=\{a_{0}\}$). Fix an aggregation correspondence $X(a_{0})$. We study the set of stochastic choice functions that are RU-rational under this restriction. Formally, a stochastic choice function $\rho$ is RU-rational given $X$ if there exist a preference distribution $\mu_{\mathscr{X}}$ and a composition distribution $\lambda$ that are consistent with the fixed aggregation correspondence $X$ such that the pair $(\mu_{\mathscr{X}},\lambda)$ rationalizes $\rho$.
As we do not make any assumptions on preferences over the underlying alternatives, the RU-rational stochastic choice functions depend only on the size of $X(a_{0})$; thus we denote the set by $RU(n)$, where $n=|X(a_{0})|$. Thus, the RU-polytope can be written as $\bigcup_{n=2}^{\infty}RU(n)$ since in the previous section we do not fix $X$.
Theorem 3.3.
Consider the outside setup (i.e., $\mathscr{A}_{N}=\{a_{0}\}$) and suppose that $\mathscr{D}=2^{\mathscr{A}}\setminus\{\emptyset\}$. Then,
$ARU\subsetneq RU(2)\subsetneq\cdots\subsetneq RU(|\mathscr{A}_{A}|)\subsetneq\underbrace{RU(|\mathscr{A}_{A}|+1)=RU(|\mathscr{A}_{A}|+2)=\cdots}_{\text{Limited\ Monotonicity + Partial RU-rationality}}.$ (6)
Theorem 3.3 shows that $RU(n)$ becomes larger as $n$ increases and when $n$ becomes larger than or equal to $|\mathscr{A}_{A}|+1$, $RU(n)$ coincides with the set of the stochastic choice functions that are characterized by Limited Monotonicity and Partial-RU rationalizability. [^16] For intuition on why the threshold is $|\mathscr{A}_{A}|+1$, see Remark D.3 in Section D.2 in the online appendix. As $n$ increases, the set $RU(n)$ becomes larger as the set of preferences over $X(a_{0})$ as well as the set of composition distributions become richer, which in turn generate richer choice behavior over $\mathscr{A}$. However, once $n$ exceeds $|\mathscr{A}_{A}|+1$, the set stabilizes and is fully characterized by the two conditions stated in Theorem 3.1. Thus
$\bigcup_{n=2}^{\infty}RU(n)=RU(|\mathscr{A}_{A}|+1)=\{\rho\text{ satisfies Limited Monotonicity \& Partial RU-rationality}\}.$
In practice, the outside option usually represents a broad heterogeneous collection of underlying alternatives—such as staying home, consuming other products, or allocating resources elsewhere—so it is reasonable to expect that the set $X(a_{0})$ is large relative to the number of explicitly modeled alternatives (i.e., $|\mathscr{A}_{A}|$). Thus, in practice, the implication of RUM can be considered to be Limited Monotonicity and Partial RU-rationality.
On the other hand, when the analyst knows $n\equiv|X(a_{0})|$ is less than $|\mathscr{A}_{A}|+1$, the sets of feasible preferences over $X(a_{0})$ and composition distributions are both smaller, rendering some behaviors not RU-rationalizable. The next remark shows that $RU(n)$ becomes non-convex: although the vertices $\rho^{\succ}_{\mathscr{E}}$ remain RU-rationalizable, some of their convex combinations do not.
@@ 3.2 Exogenously given aggregation correspondences
For any integer $n$ such that $2\leq n<|\mathscr{A}_{A}|+1$,
  • (i)
    (deterministic menu-effect) $\rho^{\succ}_{\mathscr{E}}\in RU(n)$ for all $\succ\in\mathscr{L}$ and $\mathscr{E}\subseteq\mathscr{D}$,
  • (ii)
    (non-convexity) $RU(n)$ is not convex.
Figure 1: RU polytope (blue), ARU polytope ${\mathcal{P}}_{\mathscr{A}}$ (red), and RU(3). RU(3) contains all vertices (i.e., $\rho^{\succ_{i}}_{\mathscr{E}_{i}}$) and all line segments connecting two vertices.
See Section B.3 in the online appendix for the proof. Remark 3.7 (ii) implies that $RU(n)$ admits no half-space representation; in particular, it cannot be characterized by axioms consisting solely of linear inequalities. Despite this, we can establish the following result:
@@
Fix any integer $n$ such that $2<n<|\mathscr{A}_{A}|+1$. For all $\rho$ satisfying Limited Monotonicity and Partial-RU rationality, there exists $\rho^{\prime}\in RU(n)$ such that,
$\frac{\|\rho-\rho^{\prime}\|^{2}}{|\mathscr{D}|}\leq\frac{1}{n-1}.$
The theorem implies that once $n$ is sufficiently large (e.g., $n\geq 11$), the class $RU(n)$ is highly rich and close to the set of stochastic choice functions characterized by Limited Monotonicity and Partial-RU rationality: for any $\rho$ satisfying the two axioms, there exists $\rho^{\prime}\in RU(n)$ whose average squared distance from $\rho$ is below ten percent. Therefore, when $n$ is sufficiently large, we can make the same conclusion as in the previous sections that the RU-rationality imposes much weaker restrictions than ARU-rationality.
The theorem follows by applying an approximate version of Carathéodory’s theorem. (See Theorem A.1 in the Appendix.) The key step is the following lemma:
Lemma 3.1.
For any sequence $(\succ_{i},\mathscr{E}_{i})_{i=1}^{n-1}$ with $n\geq 2$, any convex combination among $\{\rho^{\succ_{i}}_{\mathscr{E}_{i}}\}^{n-1}_{i=1}$ belongs to $RU(n)$.
See Section A.4 in the appendix for the proof of the lemma and the theorem. [^17] In Remark D.5 of Section D.2 of the online appendix, we provide an intuitive explanation why the lemma holds. Figure 1 illustrates the intuition for Theorem 3.4. The set $RU(3)$ contains all vertices (i.e., the deterministic menu-effect choice functions $\rho^{\succ_{i}}_{\mathscr{E}_{i}}$) and, moreover, every line segment connecting any two vertices. Consequently, for any RU-rational stochastic choice function $\rho$, there exists a point in $RU(3)$ arbitrarily close to $\rho$.
3.3 Multiple non-atomic aggregates
In this section, we consider the case in which there exist multiple non-atomic aggregates. The main results from the single non-atomic aggregate case continue to hold in essentially the same form, as we show below:
Theorem 3.5.
Assume Richness of $\mathscr{X}$. A stochastic choice function $\rho$ is RU-rational if and only if $\rho$ satisfies the following two conditions:
@@ 3.3 Multiple non-atomic aggregates
    (Limited Monotonicity): if $b\in D\subseteq\mathscr{A}_{A}$ and $E\subseteq\mathscr{A}_{N}$ then $\rho(D,b)\geq\rho(D\cup E,b)$,
@@
The only substantive difference from Theorem 3.1 is that the condition (i) becomes stronger. In the earlier theorem, the condition (i) required monotonicity with respect to the addition of a single aggregate—namely, the outside option $a_{0}$. In the current theorem, where multiple non-atomic aggregates are present, the condition (i) instead requires monotonicity with respect to the addition of arbitrary subsets $E$ of the non-atomic aggregates. This is a natural extension of Theorem 3.1: Theorem 3.1 corresponds to the case in which $\mathscr{A}_{N}$ is singleton $\{a_{0}\}$. Therefore, all of the implications of Theorem 3.1, as explained in the previous section hold with Theorem 3.5. We provide the formal proof in the appendix.
In addition to Theorem 3.1, all of the theorems (Theorem 3.2, 3.3, 3.4) in this section generalize to the multiple aggregated alternative case, with appropriate modifications. In Section C of the online appendix, we provide all statements and proofs of the generalized results. The vertex characterization is formalized in Theorem C.1. In the general case, the vertices of the RU-polytope are deterministic choice functions that are rationalized by a linear order $\succ$ over aggregates, with the additional behavior that the agent may arbitrarily deviate to any non-atomic aggregate. In Theorem C.2, we show that the results on exogenous aggregation correspondences also generalize in a natural way. As the number of underlying alternatives for each aggregate increase, the RU-rationalizable set expands, up until a threshold at which point Limited Monotonicity and Partial-RU characterize the set. Finally, in Theorem C.3, we show that as long as the number of underlying alternatives is relatively large, Limited Monotonicity and Partial-RU approximate the RU-rationalizable set even below the threshold. Therefore, even with multiple non-atomic aggregates, RU-rationality alone does not justify the use of ARMs.
@@ 4 Conditions restoring ARUM
This section investigates the conditions under which RU-rationality implies ARU-rationality. These conditions are especially relevant given the widespread use of ARUMs in empirical IO and the common assumption that the data-generating process satisfies RU-rationality. In particular, we provide restrictions on the underlying preference distribution $\mu_{{\scriptscriptstyle\mathscr{X}}}$ and the composition distribution $\lambda$, each of which ensures RU-rationality and ARU-rationality coincide on the observed choice domain $\mathscr{A}$. The two conditions we introduce are independent: one applies to $\mu_{{\scriptscriptstyle\mathscr{X}}}$ (Section 4.1) and the other to $\lambda$ (Section 4.2).
@@ 4.1 Condition on $\mu_{{\scriptscriptstyle\mathscr{X}}}$: non-overlappingness
In this section we provide a restriction on the distribution of preferences that recovers the implications of ARUM.
@@
Given an aggregation correspondence $X$, a distribution $\mu_{{\scriptscriptstyle\mathscr{X}}}\in\Delta(\mathscr{L}(\mathscr{X}))$ is non-overlapping if for any $\succ\in supp(\mu_{{\scriptscriptstyle\mathscr{X}}})$, $a\in\mathscr{A}_{N}$, and $y,z\in X(a)$,
@@
The condition requires that for each aggregate $a$, the underlying alternatives in $X(a)$ occupy adjacent positions in every preference ranking in the support of $\mu_{\mathscr{X}}$.  [^18] It is possible to weaken the definition as follows but we keep this condition for simplicity: A distribution $\mu_{{\scriptscriptstyle\mathscr{X}}}\in\Delta(\mathscr{L}(\mathscr{X}))$ is almost non-overlapping if for any $\succ\in supp(\mu_{{\scriptscriptstyle\mathscr{X}}})$, there exists $b,c\in\mathscr{A}$ such that (i) $y\succ x$ for all $x\in X(b)\cup X(c)$ and $y\in\mathscr{X}\setminus(X(b)\cup X(c))$ and (ii) for all $a\in\mathscr{A}_{N}\setminus\{b,c\}$, and $y,z\in X(a)$, $y\succ x\succ z\implies x\in X(a)$.
@@
Proposition 4.1 shows that, without any assumptions on $\lambda$, the use of ARUM is valid when the underlying alternatives represented by each aggregate occupy adjacent positions in agents’ preference rankings. Consequently, when constructing an aggregated dataset, alternatives with closely related characteristics should not be assigned to different aggregates. A similar observation has been made empirically by Stafford (2018) using a dataset of lobster fishing. In particular, she demonstrated that classifying crab-fishing sites as part of the outside option leads to biased coefficient estimates. Because lobster and crab fishing are close substitutes, an excellent lobster site is preferred to a poor crab site and vice-versa, their rankings overlap, thereby violating the non-overlapping condition specified in Proposition 4.1.